# Learning Nash Equilibrial Hamiltonian for Two-Player Collision-Avoiding Interactions

### v1

@@ I Introduction
Existing solutions can be categorized into learning, control, or a mixture of these two approaches: Learning approaches, e.g., variants of multi-agent reinforcement learning (MARL), approximate policies by converging both the dual (i.e., the values) and the primal (i.e., the policies) variables. However, the state-of-the-art model-free MARL, e.g., counterfactual regret minimization (CFR) [5] and fictitious self-play (FP) [6], still have difficulty to converge in games with large state/action spaces or large depth [7, 3]. Some control methods, on the other hand, find policies for given initial states using Pontryagin’s Maximum Principle (PMP) [8], and iteratively do so in real time [9, 10]. To enable fast computation, these methods leverage known system dynamics and solve depth-limited sub-games, e.g., via model predictive control. While methods for solving HJI equations exist, e.g., through level set [11, 12], these are not scalable to problems with more than a handful of state dimensions. Our method belongs to the third category, where equilibrium policies are sampled over the set of initial states and solved through PMP, and then used for learning a policy approximator in a supervised fashion [13, 14]. This approach leverages knowledge about system dynamics while generalizing solutions through neural networks.
@@ III Problem Statement
$H_{i}(u_{i};\textbf{x},\lambda_{i},t)=\lambda_{i}^{T}\boldsymbol{f}-l_{i}(\textbf{x},u_{i}),$ (1)
@@ IV-A Collision-avoiding interactions with linear dynamics
$l_{i}(\textbf{x},u_{i})=u_{i}^{2}+\phi(\textbf{x}),$ (3)
which combines an effort loss and a collision loss $\phi(\cdot)$. The latter has the following properties: There exists a non-empty subset of “collision” states $\mathcal{C}\subset\mathcal{X}\times\mathcal{X}$, such that $\phi(\textbf{x})=r$ for all $\textbf{x}\in\mathcal{C}$ and $\phi(\textbf{x})=0$ otherwise, where $r$ is a large penalty. Let $\partial\mathcal{C}$ be the boundary of $\mathcal{C}$. $\nabla_{x_{i}}\phi(\textbf{x})=0$ for all $\textbf{x}\notin\partial\mathcal{C}$, and for $\textbf{x}\in\partial\mathcal{C}$, $\nabla_{x_{i}}\phi(\textbf{x})$ is a delta function. When system dynamics is linear, i.e., $f_{i}(x_{i},u_{i})=A_{i}x_{i}+B_{i}u_{i}$, the co-state dynamics becomes
$\dot{\lambda}_{i}^{*}(t)=-A_{i}^{T}\lambda_{i}+\nabla_{x_{i}}\phi(\textbf{x}^{*}),~{}\lambda_{i}^{*}(T)=-\nabla_{x_{i}}g_{i}(x_{i}^{*}(T)).$ (4)
While $\lambda^{*}_{i}$ is in general coupled with $\textbf{x}^{*}$, Eq. (4) reveals its structure: Notice that $\lambda^{*}_{i}$ has discontinuous piece-wise linear dynamics, since for any trajectory x, $\nabla_{x_{i}(t)}\phi(\textbf{x}(t))=0$ almost everywhere on $[0,T]$ except when $x(t)\in\partial\mathcal{C}$. $\lambda^{*}_{i}$ is thus controlled by its terminal boundary $\lambda_{i}^{*}(T)$ and the set of time instances, denoted by $\mathcal{T}_{i}$, where $x(t)\in\partial\mathcal{C}$ for $t\in\mathcal{T}_{i}$. We argue that for two-player collision-avoiding interactions, we have the following five cases of $(\mathcal{T}_{1},\mathcal{T}_{2})$: (,)($\emptyset$,$\emptyset$)( ∅ , ∅ ) (no collision), ($(\{t_{1},t_{2}\},\emptyset)$ (Player 1 overtakes Player 2), (,$(\emptyset,\{t_{1},t_{2}\})$ (Player 2 overtakes Player 1), ($(\{t_{1}\},\{t_{2}\})$ (Player 1 moves toward Player 2, who then evades), and ($(\{t_{2}\},\{t_{1}\})$ (Player 2 moves toward Player 1, who then evades), where $0\leq t_{1}<t_{2}\leq T$. Sample interactions are illustrated in Fig. 1a using the uncontrolled intersection case (see below). It should be noted that while categorizing different types of collisions may seem unnecessary as collisions should be avoided in the first place, softened $\phi$, which is necessary for correctly solving BVPs, will lead to different types of non-collision close calls (i.e., when players have small minimal distances during the interaction) where $\boldsymbol{\lambda}^{*}$ has similar structures as for collisions. The equilibrium values of these different types of close calls are necessary to be learned because by telling apart these interactions one may infer the types of players [2, 4].
@@
Policy structure: With linear dynamics and Eq. (3), the Hamiltonian becomes
$H_{i}(u_{i};\textbf{x},\lambda_{i},t)=x^{T}A_{i}^{T}\lambda_{i}+u_{i}^{T}B^{T}\lambda_{i}-u_{i}^{2}-\phi(\textbf{x}).$ (5)
@@
Figure 2: Uncontrolled intersection setup with two players.
Running example: We consider a two-player uncontrolled intersection case (see Fig. 2) where each player’s state is defined by its position $d_{i}$ and velocity $v_{i}$: $x_{i}=(d_{i},v_{i})$, and action by its acceleration rate. The terminal loss is $g_{i}(x)=-\alpha d_{i}(T)+(v_{i}(T)-\bar{v})^{2}$, which encourages the agent to move forward ($\alpha>0$) and restoring a target speed $\bar{v}$ at $T$. System dynamics follows:
@@
$\dot{\lambda}_{i}^{*}[1]=\nabla_{d_{i}}\phi(\textbf{x}^{*}),~{}\lambda_{i}^{*}(T)[1]=\alpha$ (7)
@@ 18       $k=k+1$;
$a(\cdot;w)=0$ when $\hat{\textbf{y}}_{w}(\cdot)$ is fully generalizable. At the $k$-th iteration of active learning, we sample a subset $\tilde{\mathcal{X}}_{0}^{(k)}\subset\mathcal{X}_{0}$, and solve the game $\mathcal{G}(\textbf{x}^{*},t^{*})$ for $\textbf{x}^{*},t^{*}=\operatorname*{arg\,max}_{\textbf{x}\in\tilde{\mathcal{X}}_{0}^{(k)},t\in\mathcal{T}_{0}}a(\textbf{x},t;w)$. The resultant solution is then added to the dataset to update $w$. Alg. 1 summarizes the active learning algorithm, where $\mathcal{X}_{t},~{}\mathcal{X}_{a}\subset\mathcal{X}_{0}$ are the training and acquisition datasets, respectively. $\mathcal{X}_{t}\cap\mathcal{X}_{a}=\emptyset$.
@@ V-A Case setup
Using the running example on uncontrolled intersection, we learn three neural network models: one for value approximation, one for co-state approximation, and the last also for co-state approximation but through active learning. We use a variety of training data sizes for learning to observe how generalization performance improves as data increases. For each learning task, we perform twenty independent experiments to accommodate for randomness in training data sampling, model initialization, and the active learning algorithm. We consider two metrics for generalization performance: (1) the absolute prediction error in co-state at current time, and (2) the total number of collisions over a set of test cases by using the predictive model for closed-loop control. To remove confounding effects, the true equilibria of these test cases have no collisions. The second metric is of higher importance as it directly reflects the risk induced by the learned models.
@@
$\phi(x)=b\prod_{{i,j}=\{(1,2),(2,1)\}}\sigma_{1}(d_{i})\sigma_{2}(d_{j}),$
@@
$\sigma_{1}(d)=\left(1+\exp(-\gamma(d-R/2+5W/2))\right)^{-1};$ (10)
$\sigma_{2}(d)=\left(1+\exp(\gamma(d-R/2-W/2-L))\right)^{-1};$ (11)
$b=10^{4}$ sets a high loss for collision; $\gamma=5$ is a shape parameter empirically set to enable effective BVP convergence; $R$, $L$, and $W$ are the road length, car length, and car width, respectively (Fig. 2a).
Equilibrium solutions: Fig. 3a,b shows equilibrium policies and co-states sampled from $\mathcal{X}_{0}$. Consistent with our analysis, co-state trajectories have a low-dimensional representation, i.e., the trajectories over time can be compressed as four parameters. More specifically, each pair of co-state trajectories of Player 1 and 2 specifies: (1) the type of equilibrium interaction that is about to happen (see Fig. 2a) through $t_{i}^{in}$ and $t_{i}^{out}$, and (2) the risk of the interaction through $q_{i}$, where $q_{i}$ also directly governs the control policy according to Eq. (7). It is worth investigating in future study whether this representation of co-states more closely reflects what human beings learn and use to make decision during collision avoiding interactions.
@@
Learning architecture and hyperparameters: From Fig. 2b, the equilibrium value exhibits two modes along time. Therefore we propose the following value network architecture:
@@
For each learning task, we perform full batch training use a learning rate of 0.01. For early termination, we set aside 400 independently drawn validation data points, along with their BVP solutions. The training terminates when it accumulates 500 epoches where the validation error, i.e., the $l_{2}$-norm between predicted and true values or co-state trajectories, increases.
@@
Closed-loop control: For both value and co-state networks, we evaluate their generalization performance by counting their collision percentages among test cases when used for closed-loop control for both players. For the value network, we use PyTorch autograd to compute the co-state as $\nabla_{x_{i}}\hat{V}$. Following Eq. (7), the control input is then either $\hat{\lambda}_{i}[2]$ when it is within the input bounds, or otherwise one of the bounds. We roll out the simulation using a time interval of 0.03s.
@@ V-B Results
Fig. 4a compares the collision percentages among all test cases contributed by the value and the co-state networks along increasing training data sizes. We perform t-tests to show statistically significant safety improvements contributed by co-state networks. We show in Fig. 4b the mean squared errors of co-state predictions by the two networks. It is worth noting that while the performance between the two are statistically indifferent using this metric, the value network produces higher variances in its predictions under the same random procedure for sampling training data. We conjecture that the prediction variance correlates positively with the collision percentage. An inspection into this conjecture will be of interest for a future study. Fig. 5 demonstrates the test interaction trajectories along $d_{1}$ and $d_{2}$ from the ground truth, the value, and the co-state networks when the training data size is 250. Collision cases are highlighted.
@@ VI Discussion
Here we summarize a few limitations of the presented work and feasible paths towards addressing them. From linear to nonlinear dynamics: We showed that equilibrium co-states have low-dimensional embedding when the system dynamics is linear. We first note that within the scope of autonomous driving, focused interaction problems, e.g., uncontrolled intersections, roundabout, lane changing, can be approximated to have linear dynamics, in some cases under a Frenet coordinate system. In a more general scope, it is of interest to investigate whether such co-state structures can be obtained when approximated linear dynamics are learned from nonlinear systems, e.g., through Koopman operator theory [23]. From two to multiple players: The co-state structure we exploited is also contributed by the fact that there are only two players in the game, see Fig. 1a for an explanation. We conjecture that similar structures can be obtained from multi-player games when certain constraints are applied to the feasible inputs of players. For example, in an uncontrolled intersection case, the types of collisions can be enumerated, and thus the actual dimensionality of co-states can be determined, if no players can move backwards. See Fig. 1b for an example with three players.
@@ VII Conclusions
This paper is concerned about generalization performance of data-driven Nash equilibrial policies in two-player differential games. Policies of such games follow HJI, and are known to be hard to compute directly, and data-intensive when approximated through learning. We showed analytically that for collision-avoiding games with linear dynamics, the equilibrium co-states are low-dimensional due to finite equilibrium types. We then empirically showed the co-state networks lead to lower collision probabilities than value networks when trained on the same datasets. We also proposed a theory-driven active learning method where the quality of co-state prediction is evaluated by its compliance to PMP. We showed that active learning further improves collision probabilities from static learning when data size increases. Mathematically, we proposed to learn the manifold of dual variables (co-states) when it has a simple structure. Our future work will investigate the utility of this idea in multi-agent collision-avoiding interactions.

### v2

@@ I Introduction
Existing solutions can be categorized into learning, control, or a mixture of these two approaches: Learning approaches, e.g., variants of multi-agent reinforcement learning (MARL), approximate policies by converging both the dual (i.e., the values) and the primal (i.e., the policies) variables. However, the state-of-the-art model-free MARL, e.g., counterfactual regret minimization (CFR) [5] and fictitious self-play (FP) [6], still have difficulty to converge in games with large state/action spaces or large depth [7, 3]. Some control methods, on the other hand, find policies for given initial states using Pontryagin’s Maximum Principle (PMP) [8], and iteratively do so in real time [9, 10]. To enable fast computation, these methods leverage known system dynamics and solve depth-limited sub-games, e.g., via model predictive control. While methods for solving HJI equations exist, e.g., through level set [11, 12], these are not scalable to problems with more than a handful of state dimensions. Our method belongs to the third category, where equilibrium policies are sampled over the set of initial states and solved through PMP, and then used for learning a policy approximation in a supervised fashion [13, 14]. This approach leverages knowledge about system dynamics while generalizing solutions through neural networks.
@@ III Problem Statement
$H_{i}(u_{i};\textbf{x},\lambda_{i},t)=\lambda_{i}^{T}\textbf{f}-l_{i}(\textbf{x},u_{i}),$ (1)
@@ IV-A Collision-avoiding interactions with linear dynamics
$l_{i}(\textbf{x},u_{i})=u_{i}^{2}+\phi_{i}(\textbf{x}),$ (3)
which combines an effort loss and a collision loss $\phi_{i}(\cdot)$. The latter has the following properties: There exists a non-empty subset of “collision” states $\mathcal{C}\subset\mathcal{X}\times\mathcal{X}$, such that $\phi_{i}(\textbf{x})=r$ for all $\textbf{x}\in\mathcal{C}$ and $\phi_{i}(\textbf{x})=0$ otherwise, where $r$ is a large penalty. Let $\partial\mathcal{C}$ be the boundary of $\mathcal{C}$. $\nabla_{x_{i}}\phi_{i}(\textbf{x})=0$ for all $\textbf{x}\notin\partial\mathcal{C}$, and for $\textbf{x}\in\partial\mathcal{C}$, $\nabla_{x_{i}}\phi_{i}(\textbf{x})$ is a delta function. When system dynamics is linear, i.e., $f_{i}(x_{i},u_{i})=A_{i}x_{i}+B_{i}u_{i}$, the co-state dynamics becomes
$\dot{\lambda}_{i}^{*}(t)=-A_{i}^{T}\lambda_{i}+\nabla_{x_{i}}\phi_{i}(\textbf{x}^{*}),~{}\lambda_{i}^{*}(T)=-\nabla_{x_{i}}g_{i}(x_{i}^{*}(T)).$ (4)
While $\lambda^{*}_{i}$ is in general coupled with $\textbf{x}^{*}$, Eq. (4) reveals its structure: Notice that $\lambda^{*}_{i}$ has discontinuous piece-wise linear dynamics, since for any trajectory x, $\nabla_{x_{i}(t)}\phi_{i}(\textbf{x}(t))=0$ almost everywhere on $[0,T]$ except when $x(t)\in\partial\mathcal{C}$. $\lambda^{*}_{i}$ is thus controlled by its terminal boundary $\lambda_{i}^{*}(T)$ and the set of time instances, denoted by $\mathcal{T}_{i}$, where $x(t)\in\partial\mathcal{C}$ for $t\in\mathcal{T}_{i}$. We argue that for two-player collision-avoiding interactions, we have the following five cases of $(\mathcal{T}_{1},\mathcal{T}_{2})$: (,)($\emptyset$,$\emptyset$)( ∅ , ∅ ) (no collision), ($(\{t_{1},t_{2}\},\emptyset)$ (Player 1 overtakes Player 2), (,$(\emptyset,\{t_{1},t_{2}\})$ (Player 2 overtakes Player 1), ($(\{t_{1}\},\{t_{2}\})$ (Player 1 moves toward Player 2, who then evades), and ($(\{t_{2}\},\{t_{1}\})$ (Player 2 moves toward Player 1, who then evades), where $0\leq t_{1}<t_{2}\leq T$. Sample interactions are illustrated in Fig. 1a using the uncontrolled intersection case (see below). It should be noted that while categorizing different types of collisions may seem unnecessary as collisions should be avoided in the first place, softened $\phi_{i}$, which is necessary for correctly solving BVPs, will lead to different types of non-collision close calls (i.e., when players have small minimal distances during the interaction) where $\boldsymbol{\lambda}^{*}$ has similar structures as for collisions. The equilibrium values of these different types of close calls are necessary to be learned because by telling apart these interactions one may infer the types of players [2, 4].
@@
Policy structure: The system dynamics for Player $i$ and its fellow player follow $\textbf{f}=\left[\begin{array}[]{ccc}f_{i}\\ f_{-i}\\ \end{array}\right]$, where $f_{i}$ and $f_{-i}$ represent linear dynamics for both players. Using Eq. (3), the Hamiltonian becomes
$H_{i}(u_{i};\textbf{x},\lambda_{i},t)=\lambda_{i}^{T}\textbf{f}-u_{i}^{2}-\phi_{i}(\textbf{x}).$ (5)
@@
Figure 2: (a) Uncontrolled intersection setup with two players. (b) 3D visualization for collision penalty function $\phi_{i}(x)$: The dark purple region shows the actual collision area, while the red region represents a safety buffer to avoid collisions.
Running example: We consider a two-player uncontrolled intersection case (see Fig. 2a) where each player’s state is defined by its position $d_{i}$ and velocity $v_{i}$: $x_{i}=(d_{i},v_{i})$, and action by its acceleration rate. The terminal loss is $g_{i}(x)=-\alpha d_{i}(T)+(v_{i}(T)-\bar{v})^{2}$, which encourages the agent to move forward ($\alpha>0$) and restoring a target speed $\bar{v}$ at $T$. System dynamics follows:
@@
$\dot{\lambda}_{i}^{*}[1]=\nabla_{d_{i}}\phi_{i}(\textbf{x}^{*}),~{}\lambda_{i}^{*}(T)[1]=\alpha$ (7)
@@ 18       $k=k+1$;
$a(\cdot;w)=0$ when $\hat{\textbf{y}}_{w}(\cdot)$ is fully generalizable. At the $k$th iteration of active learning, we sample a subset $\tilde{\mathcal{X}}_{0}^{(k)}\subset\mathcal{X}_{0}$, and solve the game $\mathcal{G}(\textbf{x}^{*},t^{*})$ for $\textbf{x}^{*},t^{*}=\operatorname*{arg\,max}_{\textbf{x}\in\tilde{\mathcal{X}}_{0}^{(k)},t\in\mathcal{T}_{0}}a(\textbf{x},t;w)$. The resultant solution is then added to the dataset to update $w$. Alg. 1 summarizes the active learning algorithm, where $\mathcal{X}_{t},~{}\mathcal{X}_{a}\subset\mathcal{X}_{0}$ are the training and acquisition datasets, respectively. $\mathcal{X}_{t}\cap\mathcal{X}_{a}=\emptyset$.
@@ V-A Case setup
Using the running example on the uncontrolled intersection, we learn three neural network models: one for value approximation, one for co-state approximation, and the last also for co-state approximation but through active learning. We use a variety of training data sizes for learning to observe how generalization performance improves as data increases. For each learning task, we perform twenty independent experiments to accommodate for randomness in training data sampling, model initialization, and the active learning algorithm. We consider two metrics for generalization performance: (1) the absolute prediction error in co-state at current time, and (2) the total number of collisions over a set of test cases by using the predictive model for closed-loop control. To remove confounding effects, the true equilibria of these test cases have no collisions. The second metric is of higher importance as it directly reflects the risk induced by the learned models.
@@
$\phi_{i}(x)=b\sigma(d_{i},\theta)\sigma(d_{-i},1),$
@@
$\sigma(d,\theta)=$ $\left(1+\exp(-\gamma(d-R/2+\theta W/2))\right)^{-1}$ (10)
$\left(1+\exp(\gamma(d-(R+W)/2-L))\right)^{-1},$
$b=10^{4}$ sets a high loss for collision; $\gamma=5$ is a shape parameter empirically set to enable effective BVP convergence; $\theta=5$ is a parameter that sets a safety buffer zone to avoid actual collisions at the intersection. $R$, $L$, and $W$ are the road length, car length, and car width, respectively. These parameter settings and the visualization of collision penalty $\phi_{i}(x)$ are illustrated in Fig. 2.
Equilibrium solutions: Fig. 3a,b shows equilibrium policies and co-states sampled from $\mathcal{X}_{0}$. Consistent with our analysis, co-state trajectories have a low-dimensional representation, i.e., the trajectories over time can be compressed as four parameters. More specifically, each pair of co-state trajectories of Player 1 and 2 specifies: (1) the type of equilibrium interaction that is about to happen (see Fig. 2a of [2]) through $t_{i}^{in}$ and $t_{i}^{out}$, and (2) the risk of the interaction through $q_{i}$, where $q_{i}$ also directly governs the control policy according to Eq. (7). It is worth investigating in future studies whether this representation of co-states more closely reflects what human beings learn and use to make decision during collision-avoiding interactions.
@@
Learning architecture and hyperparameters: From the value landscape in Fig. 2b of [2], the equilibrium value exhibits two modes along time. Therefore we propose the following value network architecture:
@@
For each learning task, we perform full batch training use a learning rate of 0.01. For early termination, we set aside 400 independently drawn validation data points, along with their BVP solutions. The training terminates when it accumulates 500 epochs where the validation error, i.e., the $l_{2}$-norm between predicted and true values or co-state trajectories, increases.
@@
Closed-loop control: For both value and co-state networks, we evaluate their generalization performance by counting their collision percentages among test cases when used for closed-loop control for both players. For the value network, we use PyTorch autograd to compute the co-state as $\nabla_{x_{i}}\hat{V}$. Following Eq. (7), the control input is then either $0.5\hat{\lambda}_{i}[2]$ when it is within the input bounds, or otherwise one of the bounds. We roll out the simulation using a time interval of 0.03s.
@@ V-B Results
Fig. 4a compares the collision percentages among all test cases contributed by the value and the co-state networks along increasing training data sizes. We perform t-tests to show statistically significant safety improvements contributed by co-state networks. We show in Fig. 4b the mean squared errors of co-state predictions by the two networks. It is worth noting that while the performance between the two is statistically indifferent using this metric, the value network produces higher variances in its predictions under the same random procedure for sampling training data. We conjecture that the prediction variance correlates positively with the collision percentage. An inspection into this conjecture will be of interest for a future study. Fig. 5 demonstrates the test interaction trajectories along $d_{1}$ and $d_{2}$ from the ground truth, the value, and the co-state networks when the training data size is 250. Collision cases are highlighted.
@@ VI Discussion
Here we summarize a few limitations of the presented work and feasible paths towards addressing them. From linear to nonlinear dynamics: We showed that equilibrium co-states have low-dimensional embedding when the system dynamics are linear. We first note that within the scope of autonomous driving, focused interaction problems, e.g., uncontrolled intersections, roundabout, and lane changing, can be approximated to have linear dynamics, in some cases under a Frenet coordinate system. In a more general scope, it is of interest to investigate whether such co-state structures can be obtained when approximated linear dynamics are learned from nonlinear systems, e.g., through Koopman operator theory [23]. From two to multiple players: The co-state structure we exploited is also contributed by the fact that there are only two players in the game, see Fig. 1a for an explanation. We conjecture that similar structures can be obtained from multi-player games when certain constraints are applied to the feasible inputs of players. For example, in an uncontrolled intersection case, the types of collisions can be enumerated, and thus the actual dimensionality of co-states can be determined, if no players can move backwards. See Fig. 1b for an example with three players.
@@ VII Conclusions
This paper is concerned about generalization performance of data-driven Nash equilibrial policies in two-player differential games. Policies of such games follow HJI, and are known to be hard to compute directly, and data-intensive when approximated through learning. We showed analytically that for collision-avoiding games with linear dynamics, the equilibrium co-states are low-dimensional due to finite equilibrium types. We then empirically showed the co-state networks lead to lower collision probabilities than value networks when trained on the same datasets. We also proposed a theory-driven active learning method where the quality of co-state prediction is evaluated by its compliance with PMP. We showed that active learning further improves collision probabilities from static learning when data size increases. Mathematically, we proposed to learn the manifold of dual variables (co-states) when it has a simple structure. Our future work will investigate the utility of this idea in multi-agent collision-avoiding interactions.